# Towards a multi-dimensional picture of classical nucleation: The example of liquid condensation through theory and rare event sampling

### v1

@@ Abstract
While very popular for its ability to provide an elegant conceptual framework, classical nucleation theory (CNT) still fails at quantitatively reproducing most of the experimental and numerical observations of first order transitions. In this paper, we extend CNT by explicitly incorporating the multi-dimensional aspect of nucleation. Focusing on the liquid condensation of a Lennard-Jones gas, we used state of the art rare event sampling simulations to precisely characterize the nucleation event insisting on the specific features of the critical cluster. Our numerical results indicate a simultaneous growth and densification in liquid condensation. We then exploit these insights to develop a 2-variable nucleation theory. Our model based on the capillary approximation is able to quantitatively retrieve the numerical results in nucleation rate and critical cluster properties. Furthermore, our model provides a qualitatively more accurate representation of nucleation near the spinodal regime. The effectiveness of this integrated numerical and theoretical framework highlights the limitations of CNT and more recent theories while offering a robust foundation for its refinement.
@@
So far, most efforts dedicated to improving CNT consisted in revisiting the capillary model including for instance a size dependence for the surface tension [21]. Alternatively, assuming that the whole nucleation process is driven only by the nucleating cluster size has also been highly debated since the emergence of diffuse-interface models [22, 23, 24, 25]. Their predictions are in fairly good agreement with microscopic simulations [26, 27, 15], but their highly multidimensional free energy surface complicates the determination of critical cluster properties and makes the calculation of the nucleation rate very difficult. The latter has been recently attempted by Lutsko et al. for a colloidal solution [28] and by Simeone and al. within the square-gradient approximation [29]. Such fundamental frameworks are certainly very promising but their application to practical problems remains particularly challenging.
@@
With two novel ideas, our work revisits the well-studied problem of liquid condensation from vapor. On the one hand, we use an innovative numerical strategy based on rare event sampling techniques to precisely measure structural features in atomistic simulations. On the other hand, an accessible theoretical framework assuming a multi-dimensional picture of classical nucleation is built to predict structural features of the critical nucleation cluster along with the associated free energy barrier and nucleation rate. Altogether, we show that the multi-dimensional picture of classical nucleation outperfoms both qualitatively and quantitatively CNT and more recent theories, paving the way for a better understanding of nucleation.
Rare event sampling - In this first part, we will describe how different molecular dynamics (MD) simulation approaches were carried out to observe the liquid condensation from the vapor phase. In particular, we focus on the Lennard-Jones system and use the corresponding units namely energy parameter $\epsilon$, length parameter $\sigma$, Boltzmann constant $k_{\mathrm{B}}$, and monomer mass $m$. Herein, our objective is not only to simulate the liquid condensation which can be done with brute force simulations using large scale simulations at moderate supersaturation regimes [19]. Instead, we wish to isolate critical nucleation clusters and generate a representative ensemble of structures located near the saddle-point of the free energy barrier, enabling a precise characterization of their properties. Depending on the supersaturation regime, two different approaches are carried out.
In the lowest supersaturation regimes (the vapor density $\rho_{0}<0.015\sigma^{-3}$), the cluster size is expected to be large and we therefore are able to use the seeding technique [43] yet with a particular implementation that allows us to stabilize critical clusters. For the initialization, a spherical liquid droplet is artificially inserted into a gas phase. When the density and the box size are correctly chosen, the droplet spontaneously stabilizes at a particular radius and density which are most probably different from the initial conditions. Meanwhile, the external gas density is then automatically adjusted through finite-size effects and mass conservation at a new equilibrium value $\rho_{0}$ for which the obtained liquid droplet is a critical nucleation cluster [44]. This behavior is known as ”super-stabilization” effect in confined systems [45, 46, 47, 48, 15, 49]: Ultimately, by testing different box sizes and average densities, one can stabilize liquid clusters at different supersaturation regimes. Then, an ensemble of critical clusters can be obtained upon changing the initial velocity and atomic positions.
@@
Theory - Our theoretical framework builds on the multi-variable nucleation theory [55, 56]. To simplify its application, we extend the capillary approximation in CNT by explicitly incorporating both the size $R$ and the density $\rho$ of the nucleating cluster. Consequently, the work of formation is given by $\Delta\varOmega(R,\rho)=-4\mathrm{\pi}{R}^{3}g_{n}/3+4\mathrm{\pi}{R}^{2}\gamma$. Here, $g_{n}$ is the driving force for nucleation, defined as $g_{n}=\mathcal{F}(\rho_{0})+\frac{\partial\mathcal{F}}{\partial\rho}\Big{|}_{\rho=\rho_{0}}(\rho-\rho_{0})-\mathcal{F}(\rho)$, with $\mathcal{F}$ the Helmholtz free energy per unit volume [23]. The surface tension of vapor-liquid interface, $\gamma$, is inspired by the Square Gradient Approximation (SGA), $\gamma=\kappa(\rho-\rho_{0})^{2}$, where $\kappa$ is assumed density-independent and can be computed from the surface tension at equilibrium [47, 57, 58, 25, 59]. We note that unlike in CNT, both $g_{n}$, and $\gamma$ are now treated as functions of the cluster density, $\rho$.
@@
The nucleation trajectory passes through the saddle point of the work of formation, $(R_{\mathrm{c}},\rho_{\mathrm{c}})$, defining the critical cluster size and density. The corresponding work of formation is $\Delta\varOmega_{\mathrm{c}}$, and the Hessian matrix at the critical point is $\mathbf{H}$. Near this saddle point,
@@
Firstly, for the distribution parameter $f_{0}$, the assumption made by Alekseechkin [56] regarding the diagonal elements of $\mathbf{H}$ does not always hold. Instead, following Langer’s idea [55], we define a line passing through the saddle point, perpendicular to the nucleation flux, and separating metastable and stable regions. Nuclei exist only in the metastable region (MS), allowing $f_{0}$ to be determined via mass conservation:
@@
where $q(R,\rho)=4\mathrm{\pi}\rho R^{3}/3$ is the number of monomers in a cluster.
Secondly, the diffusivity matrix $\mathbf{D}$ is derived from the formulation of dynamical density functional theory (DDFT) [61, 62, 63] near the critical cluster. DDFT is based on the continuity equation with the assumption that the driving force for particle diffusion is proportional to the real-space diffusion constant $D_{\mathrm{diff}}$ and the gradient of chemical potential [64]. A detailed explanation can be found in Ref. [65] and Appendix. In practice, the diffusion constant is estimated from our seeding MD simulations of the early growth stages of post-critical clusters [66]. At $T=0.8\epsilon/k_{\mathrm{B}}$, we find $D_{\mathrm{diff}}=0.03\sigma\sqrt{\epsilon/m}$.
@@
Finally, surface energy factor at this temperature is determined from the surface tension at equilibrium ($\gamma^{\mathrm{eq}}=0.8633\epsilon/\sigma^{2}$) [75] using $\kappa=\gamma^{\mathrm{eq}}/(\rho^{\mathrm{eq}}_{1}-\rho^{\mathrm{eq}}_{2})^{2}$.
@@
Results - To begin, from Fig. 1(a) which shows representative critical clusters obtained with our rare-event sampling strategy, one can already observe that along with the size, the density of the nucleating cluster seem to decrease when increasing supersaturation. Then, more quantitatively, Fig.1 (b-c) show the critical cluster radii and densities averaged over the whole distribution of clusters obtained in simulations. As predicted by other multi-dimensional nucleation [76, 47, 57], our numerical results confirm that the density of the critical cluster decreases with supersaturation. From the theoretical viewpoint, on the one hand, the critical cluster size is similarly predicted by all models. At the lowest supersaturation, there is a strong agreement with the simulation data. However, all models slightly underestimate the cluster size at intermediate supersaturation levels. On the other hand, regarding the critical cluster density, CNT fails at capturing the observed densification. Meanwhile, although it is able to qualitatively retrieve this decrease, the SGA diffuse-interface model [23, 60, 27, 29] overestimates the cluster density and exhibits a maximum that is not seen in simulations. In contrast, our two-variable model successfully reproduces the numerical results both qualitatively and quantitatively.
@@
Figure 1: (a) Snapshots of critical clusters from our simulations and evolution of the (b) radius and (c) density of critical cluster against initial density at temperature $T=0.8\epsilon/k_{\mathrm{B}}$. The first three snapshots are from seeding simulation, and the last two are from metadyanmics simulation. Our seeding and aimless shooting MD simulation results are represented by blue squares and circles, respectively, with error bars indicating standard deviations. Predictions from our two-variable nucleation theory, CNT, and SGA diffuse-interface model are represented by red solid lines, yellow dotted lines, and black dash lines, respectively.
Next, we use our theoretical model and CNT to evaluate the steady-state nucleation rate [See Fig.2]. It is important to note that computing the nucleation rate within the SGA model is very difficult [29, 28] and not intended here. Then, we confront those predictions to numerical results obtained in the literature or using our own brute force simulations reported in Table 1 in Appendix and analyzed with the Yasuoka-Matsumoto (YM) threshold method [33, 19]. Remarkably, the two-variable model performs very well at high supersaturation, where CNT is known to fail by several orders of magnitude. Our model contains no parameters adjusted to nucleation simulations, which demonstrates its predictive capability when properly applied and coupled to quantitative thermodynamics and kinetics data. Furthermore, contrary to CNT, the nucleation rate shows a maximum near the spinodal, an effect also observed in diffuse-interface models [29].
@@ Appendix B Simulation analysis
Cluster radius $R$, density $\rho$, and initial density $\rho_{0}$ from simulations were determined by fitting the integral of a modified sigmoid function [86],
@@ Appendix C Quantitative outputs from the simulation
$0.035$ $0.556\pm 0.052$ $3.28\pm 0.29$ $1.2\times 10^{-8}$
$0.04$ $0.490\pm 0.065$ $3.36\pm 0.17$ $4.7\times 10^{-8}$
$0.05$ $3.1\times 10^{-7}$
$0.06$ $2.3\times 10^{-6}$
$0.07$ $6.1\times 10^{-6}$
$0.08$ $6.8\times 10^{-6}$

### v2

# Towards a multi-dimensional picture of nucleation: The example of liquid condensation through theory and rare event sampling

@@ Abstract
While very popular for its ability to provide an elegant conceptual framework, classical nucleation theory (CNT) often fails at quantitatively reproducing most of the experimental and numerical observations of first order transitions. In this paper, we extend CNT by explicitly incorporating the multi-dimensional aspect of nucleation. Focusing on the liquid condensation of a Lennard-Jones gas, we used state of the art rare event sampling simulations to precisely characterize the nucleation event emphasizing the specific features of the critical cluster. Our numerical results indicate a simultaneous growth and densification in liquid condensation. We then exploit these insights to develop a 2-variable nucleation theory. Our model based on the capillary approximation is able to quantitatively retrieve the numerical results in nucleation rate and critical cluster properties. Furthermore, our model provides a qualitatively more accurate representation of nucleation near the spinodal regime. The effectiveness of this integrated numerical and theoretical framework highlights the limitations of CNT and more recent theories while offering a robust foundation for its refinement.
@@
So far, most efforts dedicated to improving CNT consisted in revisiting the capillary model including for instance a size dependence for the surface tension [21]. Alternatively, assuming that the whole nucleation process is driven only by the nucleating cluster size has also been highly debated since the emergence of diffuse-interface models [22, 23, 24, 25]. Their predictions are in fairly good agreement with microscopic simulations [26, 27, 15], but their highly multidimensional free energy surface complicates the determination of critical cluster properties and makes the calculation of the nucleation rate very difficult. The latter has been recently attempted by Lutsko et al. for a colloidal solution [28] and by Simeone et al. within the square-gradient approximation [29]. Such fundamental frameworks are certainly very promising but their application to practical problems remains particularly challenging.
@@
With two novel ideas, our work revisits the well-studied problem of liquid condensation from vapor. On the one hand, we use an innovative numerical strategy based on rare event sampling techniques to precisely measure structural features in atomistic simulations. On the other hand, an accessible theoretical framework assuming a multi-dimensional picture of nucleation is built to predict structural features of the critical nucleation cluster along with the associated free energy barrier and nucleation rate. Altogether, we show that the multi-dimensional picture of nucleation outperforms both qualitatively and quantitatively CNT and more recent theories, paving the way for a better understanding of nucleation.
Rare event sampling - In this first part, we will describe how different molecular dynamics (MD) simulation approaches were carried out to observe the liquid condensation from the vapor phase. In particular, we focus on the Lennard-Jones system and use the corresponding units namely energy parameter $\epsilon$, length parameter $\sigma$, Boltzmann constant $k_{\mathrm{B}}$, and monomer mass $m$. Herein, our objective is not only to simulate the liquid condensation which can be done with brute-force simulations using large scale simulations at moderate supersaturation regimes [19]. Instead, we wish to isolate critical nucleation clusters and generate a representative ensemble of structures located near the saddle point of the free energy barrier, enabling a precise characterization of their properties. Depending on the supersaturation regime, two different approaches are carried out.
In the lowest supersaturation regimes (the vapor density $\rho_{0}<0.015\sigma^{-3}$), the cluster size is expected to be large and we therefore are able to use the seeding technique [43] yet with a particular implementation that allows us to stabilize critical clusters. For the initialization, a spherical liquid droplet is artificially inserted into a gas phase. When the density and the box size are correctly chosen, the droplet spontaneously stabilizes at a particular radius and density which are most probably different from the initial conditions. Meanwhile, the external gas density is then automatically adjusted through finite-size effects and mass conservation at a new equilibrium value $\rho_{0}$ for which the obtained liquid droplet is a critical nucleation cluster [44]. This behavior is known as ”super-stabilization” effect in confined systems [45, 46, 47, 48, 15, 49]. Ultimately, by testing different box sizes and average densities, one can stabilize liquid clusters at different supersaturation regimes. Then, an ensemble of critical clusters can be obtained upon changing the initial velocity and atomic positions.
@@
Theory - Our theoretical framework builds on the multi-variable nucleation theory [55, 56]. To simplify its application, we extend the capillary approximation in CNT by explicitly incorporating both the size $R$ and the density $\rho$ of the nucleating cluster. Consequently, the work of formation is given by $\Delta\varOmega(R,\rho)=-4\mathrm{\pi}{R}^{3}g_{n}/3+4\mathrm{\pi}{R}^{2}\gamma$. Here, $g_{n}$ is the driving force for nucleation, defined as $g_{n}=\mathcal{F}(\rho_{0})+\frac{\partial\mathcal{F}}{\partial\rho}\Big{|}_{\rho=\rho_{0}}(\rho-\rho_{0})-\mathcal{F}(\rho)$, with $\mathcal{F}$ the Helmholtz free energy per unit volume [23]. The surface tension of vapor-liquid interface, $\gamma$, is inspired by the Square Gradient Approximation (SGA), $\gamma=\kappa(\rho-\rho_{0})^{2}$, where $\kappa$ is assumed density-independent and can be computed from the surface tension at equilibrium [47, 57, 58, 25, 59]. For simplification purpose, we chose to ignore the dependence of surface tension with size [21]. We note that unlike in CNT, both $g_{n}$, and $\gamma$ are now treated as functions of the cluster density, $\rho$.
@@
The nucleation trajectory passes through the saddle point of the work of formation, $(R_{\mathrm{c}},\rho_{\mathrm{c}})$, defining the critical cluster size and density. The corresponding work of formation is $\Delta\varOmega_{\mathrm{c}}$, and the Hessian matrix at the saddle point is $\mathbf{H}$. Near this saddle point,
@@
Firstly, for the distribution parameter $f_{0}$, neither Alekseechkin’s nor Langer’s analytical expression is applicable to our studies: Alekseechkin’s assumption regarding the diagonal elements of $\mathbf{H}$ [56] is not universally valid, and Langer’s assumption regarding the metastable minimum is sharp and well isolated [55] does not hold at all in our two-variable framework. We propose to employ mass conservation in the metastable region (MS) to calculate $f_{0}$,
@@
where $q(R,\rho)=4\mathrm{\pi}\rho R^{3}/3$ is the number of monomers in a cluster. We emphasize here that the determination of $f_{0}$ is extremely difficult in high dimensional phase space, which makes the calculation of nucleation rate hardly achieved, while our two-dimensional framework and calculating method of $f_{0}$ provide a simple way to obtain nucleation rate. Unlike in Ref. [59], our calculating method of $f_{0}$ is normalizable.
Secondly, the diffusivity matrix $\mathbf{D}$ is derived from the formulation of dynamical density functional theory (DDFT) [61, 62, 63] near the saddle point. DDFT is based on the continuity equation with the assumption that the driving force for particle diffusion is proportional to the real-space diffusion constant $D_{\mathrm{diff}}$ and the gradient of chemical potential [64]. Thus, mass conservation is locally considered for the growth of clusters near the saddle point. A detailed explanation for the calculation of $\mathbf{D}$ can be found in Ref. [65] and Appendix. In practice, the diffusion constant is estimated from our seeding MD simulations of the early growth stages of post-critical clusters [66]. At $T=0.8\epsilon/k_{\mathrm{B}}$, we find $D_{\mathrm{diff}}=0.03\sigma\sqrt{\epsilon/m}$.
@@
Finally, surface energy factor at this temperature is determined from the surface tension at equilibrium ($\gamma^{\mathrm{eq}}=0.8633\epsilon/\sigma^{2}$[75] using $\kappa=\gamma^{\mathrm{eq}}/(\rho^{\mathrm{eq}}_{1}-\rho^{\mathrm{eq}}_{2})^{2}$.
@@
Results - To begin, from Fig. 1(a) which shows representative critical clusters obtained with our rare-event sampling strategy, one can already observe that along with the size, the density of the nucleating cluster seem to decrease when increasing supersaturation. Then, more quantitatively, Fig. 1 (b-c) show the critical cluster radii and densities averaged over the whole distribution of clusters obtained in simulations. As predicted by other multi-dimensional nucleation theories [76, 47, 57], our numerical results confirm that the density of the critical cluster decreases with supersaturation. From the theoretical viewpoint, on the one hand, the critical cluster size is similarly predicted by all models. At the lowest supersaturation, there is a strong agreement with the simulation data. However, all models slightly underestimate the cluster size at intermediate supersaturation levels. On the other hand, regarding the critical cluster density, CNT fails at capturing the observed densification. Meanwhile, although it is able to qualitatively retrieve this decrease, the SGA diffuse-interface model [23, 60, 27, 29] overestimates the cluster density and exhibits a maximum that is not seen in simulations. In contrast, our two-variable model successfully reproduces the numerical results both qualitatively and quantitatively.
@@
Figure 1: (a) Snapshots of critical clusters from our simulations and evolution of the (b) radius and (c) density of critical cluster against initial density at temperature $T=0.8\epsilon/k_{\mathrm{B}}$. The first three snapshots are from seeding simulation, and the last two are from steering + aimless shooting simulation. Our seeding and aimless shooting MD simulation results are represented by blue squares and circles, respectively, with error bars indicating standard deviations. Predictions from our two-variable nucleation theory, CNT, and SGA diffuse-interface model are represented by red solid lines, yellow dotted lines, and black dash lines, respectively.
Next, we use our theoretical model and CNT to evaluate the steady-state nucleation rate [See Fig. 2]. Computing the nucleation rate within the SGA model and comparing it with our two-variable theory would be very valuable since both approaches go beyond the simplistic one-variable CNT. However, this calculation is very difficult [29, 28] and not performed here. Then, we compare those predictions to numerical results obtained in the literature or using our own brute-force simulations reported in Table 1 in Appendix and analyzed with the Yasuoka-Matsumoto (YM) threshold method [33, 19]. Remarkably, the two-variable model performs very well at high supersaturation, where CNT is known to fail by several orders of magnitude. Our model contains no parameters adjusted to nucleation simulations, which demonstrates its predictive capability when properly applied and coupled to quantitative thermodynamics and kinetics data. Furthermore, contrary to CNT, the nucleation rate shows a maximum near the spinodal, an effect also observed in diffuse-interface models [29].
@@ Appendix A Simulation protocol
At high supersaturation, brute-force simulations were performed in the NVT ensemble, in boxes of size $L=100\sigma$. 5 independent simulations were launched for each initial density. This change in system sizes allows us to obtain many nucleation events at the same time thus matching with the Yasuoka-Matsumoto method.
@@ Appendix B Simulation analysis
Cluster radius $R$, density $\rho$, and vapor density $\rho_{0}$ from simulations were determined by fitting the integral of a modified sigmoid function [86],
@@ Appendix C Quantitative outputs from the simulation
$0.035$ $0.556\pm 0.052$ $3.28\pm 0.29$ $1.3\pm 0.3\times 10^{-8}$
$0.04$ $0.490\pm 0.065$ $3.36\pm 0.17$ $9.1\pm 0.8\times 10^{-8}$
$0.05$ $4.9\pm 1.2\times 10^{-7}$
$0.06$ $1.7\pm 0.3\times 10^{-6}$
$0.07$ $3.0\pm 0.4\times 10^{-6}$
$0.08$ $4.3\pm 0.7\times 10^{-6}$